# What’s DAT? Three Case Studies of Measuring Software Development Productivity at Meta With Diff Authoring Time

### v1

@@ 3.3. Additional DAT Heuristic: Anchor sessions
DAT includes additional heuristics beyond precise matches. These were developed to capture more of the time surrounding diff creation, especially in tools not supported by DAT telemetry. The idea generally is to map coding activity preceding a precise match to the same diff as the precise match, recognizing that related activities often occur sequentially.
@@
For instance, in Figure 1, terminal work cannot be precisely matched due to lack of Sapling integration, but Anchor-DAT maps it to D123 because it precedes a precise match. This conservative method captures anchor sessions leading up to a diff, which are otherwise hard to trace. Activity is captured only for a limited time before a precise match and only if it involves coding-related tools. The definition of coding-related tool is shared with TSD, thus duration covered by DAT is a true subset of the duration covered by TSD.
4. Validation
4.1. User Experience Research Study
@@
Our first step in validating DAT was to create a ground truth (GT) dataset of how long developers spent on a diff. This allowed us to assess our current DAT version’s accuracy by comparing it with GT data and observing trends as we improved the algorithms.
We conducted a User Experience Research (UXR) study, screen-recording volunteer developers working on a diff. To counteract biases from screen-recording and volunteering, we used a stratified random selection of developers and diffs for broad coverage. A UX researcher transcribed video actions, noting tool use and relevance to diff development. In the future, this could train an AI to derive ground truth timings from video data.
We analyzed over a dozen developers’ diffs, achieving an average accuracy of over 90%, meaning the computed DAT was on average within ±$\pm 5\%$ of its GT value. In later DAT versions, the algorithm’s accuracy often surpassed human labelling, as developers rapidly switched between diffs, challenging human annotators.
4.2. Large Scale Survey
@@
Figure 4. Human DAT estimates (solid) and actual DAT values for an earlier DAT version (dashed).
While the UXR study was crucial for testing and development, it was limited in scale. Moreover, acceptance of a metric depends on how well it matches experience. Therefore, we conducted a survey asking developers about the time spent on a diff. This survey was integrated into the Phabricator tool, appearing once a diff was shipped to ensure responders had the full development context fresh in their minds.
We received responses for 968 unique diffs to compare with DAT values. As Figure 4 shows, DAT and survey data generally align, with developer-reported times being slightly larger, a trend also seen in the UXR study (Section 4.1). The difference between survey and DAT values increases for longer diffs, but this drift has been mitigated with newer DAT versions.
4.3. Descriptive Statistics
We performed two main validations for DAT at an aggregate level—diff coverage and diff duration. These validations were intended to verify that DAT covers the expected number of diffs and that the overall time spent in these diffs yields reasonable values. As a sanity check, the coverage and duration validations were done by comparing the current DAT version to its predecessor. The duration validations also included comparison between user-level DAT and TSD (cf. Section 2.4).
We performed several tests based on our expectations of how much additional heuristics should increase DAT. First, we expected anchor sessions durations to be higher than precise match durations. Second, we expected that the sum of all precise matches and anchor session durations never be larger than the sum of all TSD for a user during a particular time window. Any exceptions would only be possible if there is a mistake in the data or our algorithms, for example double-counting time. A final demand on any top-line metric is that it be stable over time—Figure 3 shows this.
Before releasing DAT, we ensured it passed all these checks. We found that DAT covers 87% of all eligible diffs. DAT’s duration has a right-skewed distribution, which means a small number of extreme values can significantly impact the mean. To account for slower diffs while still being sensitive to changes in the distribution, we use the 99th percentile winsorized mean to report on Anchor-DAT, which was very plausible with regard to other validations.
4.4. Time-Line Visualization
@@
Figure 5 visualizes an example diff’s DAT and how raw telemetry is covered by an additional heuristic, Anchor-DAT. DAT uses raw telemetry sessions, including start and stop times of coding tools, shown in the bottom row and maps each session to a diff, if possible. The raw sessions already include noise-removal, like pauses for inactivity (e.g., around 12:45). During DAT additional heuristic development, we also cross-validated with authors to ensure extrapolations were sensible. These checks helped refine our understanding, re-evaluate heuristic assumptions, and adjust parameters like time thresholds.
@@ 5.1. Context
Hack is one of our server-side languages. Originating from PHP, it combines a JIT-based runtime (Ottoni, 2018) with a gradual, object-oriented type system (providing static correctness guarantees similar to TypeScript for JavaScript). Hack developers can depend on commonly shared frameworks, including mocking for testing (Spadini et al., 2017). While CI forces test execution for every diff, developers can additionally run them locally if they so wish.
Hack code is statically typed to varying degrees. In the presence of suitable type annotations, Hack has the ability to detect certain type errors through static analysis, which are reported to developers directly in VS Code. In addition to static analysis, the Hack runtime enforces the arguments passed to functions, and ensures that they match the declared parameter types, throwing a dynamic TypehintViolation-exception in case of a mismatch. In particular, the standard mocking framework used to expose an untyped API, using the catch-all type mixed in Listing 6(b). As a result, its use did not enjoy any static guarantees. Values flowing through ‘mixed‘ could encounter runtime enforcement boundaries, potentially leading to runtime crashes.
@@ 5.2. Problem Statement
In the example diff in Listing 6(a), a developer improves the Application to be more object-oriented, introducing a new UserData class, and altering existing code (lines 13 and 19) to make efficient use of it. However, the developer might not be aware of the existing test.
In the case of the old, untyped mocking framework (Listing 6(b)), they would learn that they had broken a test only after CI ran, introducing significant delays: since the mocked method-body now returns a dict value where UserData is expected, the Hack runtime throws a TypehintViolation. This leads to an additional cycle of checking out a diff, reproducing and debugging the unexpected failure, coming up with the right fix, and verifying it. The lack of static types for mocking has proven to be a hindrance to productivity, as indicated by the disproportionately high number of TypehintViolation-exceptions from CI logs.
@@
5.3. Experimental Setup
Our experiments showed that we would be able to automatically migrate the majority (82%) of calls into the untyped mocking API onto the new typed variant (migratable call sites). The remainder of call sites (or unmigratable call sites) exercised the mocking API in unconventional ways and could not be auto-migrated.
We proceeded to first migrate a random selection of half of the migratable call sites ($\sim$169,000 or 42% of all call sites in the codebase). During a period of three weeks, we categorised all committed diffs:
  • Control Group (53.1% of diffs): Diffs that modified at least one migrated, but no unmigrated file.
  • Test Group (24.5% of diffs): Diffs that modified at least one unmigrated, but no migrated files.
  • Mixed Group (22.4% of diffs): Diffs that modified at least one migrated and one unmigrated file.
To estimate the effect of typed mocking, we employ a two-tailed Welch’s t-test, testing on the mean DAT difference between control and test groups. We discarded the mixed group as it contained diffs that touched both migrated and unmigrated files. Because bigger diffs naturally require more coding time, we controlled for the diff size confounder using stratification.
@@ 5.4. Results
# of files changed % DAT saved $\frac{|Control|}{|Test|}$ $p$
1 $\sim 22\%$ $2.5$ $<0.001$
2 $\sim 22\%$ $2.3$ $<0.001$
@@
Variable $\sim 14\%$ $2.2$ $<0.001$
Table 1 summarizes the results of the stratified t-tests, revealing that use of the typed mocking API significantly reduced the time to write diffs, even after controlling for the number of files in each diff. The table shows the percentage of DAT saved, the relative control and test group sizes, and the p-value for each diff size category; diff size is measured by the number of files changed in a diff.
For diff sizes up to 4 files, the typed mocking API resulted in an approximate 19–24% reduction in DAT, with all diff size categories showing statistically significant reductions. When including all diffs regardless of diff size, we measure a statistically significant reduction of approximately $14\%$ DAT ($p<0.001$). This is lower than the values in Table 1, as the inclusion of larger diffs (with more than four files changed), which only tangentially touch mocking-related files, dilutes the real savings seen in lower diff size strata.
To the best of our knowledge, this is the first time a controlled, large-scale experiment demonstrated the productivity benefits of typing in industry. While the results of this experiment have helped to quantify the direct effect of a Hack language feature and its use on the company’s productivity, the wider organizational decision-making processes also benefit. Moreover, this result demonstrates that investing in low-latency static analysis tooling to shift correctness signals early in the development cycle can lead to substantial improvements in developer productivity, especially in large code bases.
@@ 6.1. Context
React is Meta’s JavaScript framework for developing dynamic user interfaces (Gackenheimer, 2015). Since its open-source release in 2013, it has become a standard for web and mobile apps (Facebook Inc., 2013). React components generate user interfaces based on the application’s current state. When the state changes, React updates the interface by re-evaluating components. To optimize performance, developers must use memoization to prevent recomputing unchanged components.
@@ 6.2. Problem Statement
React developers can memoize in two ways:
  1. (1)
    Manual memoization, which involves writing code to save component states manually and for every component independently.
  2. (2)
    Auto-memoization, available in recent React versions, uses a new Meta-built compiler feature to automatically memoize components, representing a significant investment by the React team.
Manual memoization is thought to be time-consuming, prone to bugs, and does not scale, but the magintude of its impact on productivity remains unknown.
@@ 6.3. Methodology and Results
Figure 7. Auto-Memoization Case Study results show beanplots of DAT for manual (left) and automatic memoization (right) in a small subset of React diffs.
To understand what difference the new React compiler might make to development velocity, we first compared diffs with auto-memoization to those without. We find that the DAT median, quartiles, and skew are significantly higher for diffs with manual memoization.
To account for confounding variables such as lines of code, developer tenure, organization, and time (modelled as month of the year), we built a mixed effect regression model. We computed %$\Delta\mbox{DAT}$ to measure differences in mean DAT distributions, shown in Figure 7. For non-randomized data, we used the Wasserstein distance to measure the true difference between groups. This metric provides the minimal transport distance between two probability distributions (Wikipedia, 2024). It is robust to noise and outliers, making it ideal for comparing distributions with different shapes or scales. The final results put %$\Delta\mbox{DAT}$ between diffs with manual memoization and without memoization at 33%.
@@ 7.1. Context
Meta’s development frameworks enable cross-platform code sharing, allowing code to be written once and deployed across multiple apps and platforms, reducing redundant efforts. One well-known example of such a framework is the open-sourced React framework (Facebook Inc., 2013). Using such frameworks supposedly saves time compared to writing platform-specific code, because it only needs to be written once, but can be deployed to a number of different platforms. Developers report feeling more efficient with these frameworks, yet quantifying these gains before DAT was impossible.
@@ 7.2. Methodology
Counterfactual DAT aims to provide a realistic estimation for how much DAT diffs would have taken if code sharing was not used. While DAT for the code-shared diffs is easy to query, we cannot directly measure counterfactual development time without sharing because these hypothetical diffs were never written. Instead, we estimate it by identifying the platforms (such as Android) and apps (such as Instagram or Threads) targeted in each code-shared diff. To obtain a baseline for each platform-app combination, we compute a trimmed mean DAT for unshared diffs. We normalize by diff size, dividing diffs into terciles based on modified lines of code.
The total counterfactual DAT is then the sum of all targeted platform-app combinations in the code-shared diff, controlled for diff size. A consequence of this is that $DAT_{saved}$ could be negative, in case a code-shared diff incurs overhead possibly due to the complex nature of shared libraries. Conversely, the higher the platform-app count is, the greater the potential for actual savings.

### v2

@@ 3.2.2. Additional DAT Heuristic: Anchor-DAT
DAT offers a heuristic beyond precise matches. It was developed to easily capture more of the time surrounding diff creation, especially in tools that have no notion of diffs or commits, such as interactive playgrounds—e.g., a SQL interface. The idea is to map coding activity preceding a precise match to the same diff as the precise match, recognizing that related activities mostly co-occur sequentially.
@@
For instance, in Figure 1, work in the external terminal application cannot be precisely matched due to lack of Sapling integration, but additional heuristics map it to D123 because it precedes a precise match VS Code session to D123. This conservative method captures anchor sessions leading up to a diff, which are otherwise hard to trace. Activity is captured only for a limited time before a precise match (a few minutes) and only if it involves a list of manually vetted coding-related tools. This list is shared with TSD, making DAT a subset of the duration covered by TSD. Whenever we refer to DAT in the following, we refer to the combination of precise matches and Anchor-DAT.
3.3. Validation
3.3.1. User Experience Research Study
@@
Our first step in validating DAT was to create a ground truth (GT) dataset of how much time developers spend on a diff. This allowed us to assess DAT’s accuracy by comparing it with GT data and observing trends as we improved the algorithms and tweaked its main parameters (namely, the list of to-be-supported internal coding tools, the anchor length, and inactivity timeout).
With a team of User Experience (UX) researchers we conducted a study in which we screen-recorded developers working on a diff from inception to landing, then transcribed their actions and the duration of them. Finally, we compared the results to the extracted DAT sessions on a by-second basis. To counteract biases from screen-recording and volunteering, we used a stratified random selection of developers and diffs for broad coverage, ensuring that different strata were present in our sample (e.g., small diffs, stacked diffs, different languages and IDEs, different developer roles).
We analyzed $n=\sim 20$ diffs, achieving an average accuracy of over 90%, meaning the computed DAT was on average within $\pm 5\%$ of its GT value. In later iterations of the algorithm, DAT’s accuracy surpassed human labeling for many of the analyzed diffs, as developers rapidly switched between tools, activities, and diffs. These sometimes sub-second context switches made it hard for the UX annotators to keep track of developers’ work items at all times, especially when multiple diffs were developed in parallel. When we revisited the source video for the diverging GT transcriptions, we often found DAT to be correct.
3.3.2. Large Scale Survey
Figure 3. Survey-estimated (solid) vs. actual DAT values show the small systematic lower bound of DAT compared to human estimates.
@@ 3.3.2. Large Scale Survey
We found that acceptance of an important top-line metric depends on how well it matches experience and intuition. Therefore, we conducted a large-scale survey asking developers to estimate the time spent on a diff. This one-question survey was integrated into the Phabricator tool, appearing once a diff was shipped to ensure responders had the full development context fresh in their minds.
We received responses for 968 unique diffs to compare with DAT values. As Figure 3 shows, DAT and survey data generally align, with developer-reported times being slightly higher, a trend also seen in the UX study (Section 3.3.1). We observe that DAT remains a floor for all measured DAT percentiles.
3.3.3. Descriptive Statistics
We performed two main validations for DAT at an aggregate level—diff coverage and diff duration. These validations were intended to verify that DAT covers the expected number of diffs (diff coverage) and that the overall time spent in these diffs yields reasonable values. As a sanity check, the coverage and duration validations were done to surpass pre-defined goals. The duration validations also included comparison between the summed-up DAT for a given time period and user and their respective TSD (cf. Section 2.2), where $\sum DAT\leq TSD$
We performed several tests based on our expectation of how much Anchor-DAT should increase the precise matches. First, we expected anchor sessions’ durations to be higher than precise match durations. Second, we expected that the sum of all precise matches and anchor session durations never be larger than the sum of all TSD for a user during a particular time window. Any exceptions would only be possible if there is a mistake in the data or our algorithms, for example due to double-counting time. A final demand on a top-line metric is to be stable over time—Figure 4 shows this.
Before releasing DAT, we ensured it passed all these checks. We found that DAT covers ¿95% of all eligible diffs. DAT’s duration has a right-skewed distribution, which means a small number of extreme values can significantly impact the mean. To account for longer diffs while still being sensitive to changes in the distribution, we use the 99th percentile winsorized mean to report on Anchor-DAT, which yielded plausible results consistent with the other validations.
3.3.4. Time-Line Visualization
@@
Figure 5 visualizes raw telemetry (bottom row) from a two hour time frame, how it is mapped to an example diff’s precise DAT (top row), and how auxiliary tools are covered via Anchor-DAT. DAT uses raw telemetry sessions, including start and stop times of coding tools, shown in the bottom row. DAT tries to map each session to a diff. The raw telemetry sessions already include noise-removal, such as pauses for inactivity (e.g., around 12:45). During Anchor-DAT development, we also cross-validated with authors to ensure extrapolations were sensible. These checks helped refine our understanding, re-evaluate heuristic assumptions, and adjust parameters like time thresholds.
@@ 4.1. Context
Hack is one of Meta’s server-side languages. Originating from PHP, it combines a JIT-based runtime (Ottoni, 2018) with a gradual, object-oriented type system, providing static correctness guarantees similar to TypeScript for JavaScript. Hack developers can depend on commonly shared frameworks, including mocking for testing (Spadini et al., 2017). While CI forces test execution for every diff, developers can additionally run them locally if they so wish.
Hack code is statically typed to varying degrees. In the presence of suitable type annotations, Hack has the ability to detect certain type errors through static analysis, which are reported to developers directly in VS Code. In addition to static analysis, the Hack runtime enforces the arguments passed to functions, and ensures that they match the declared parameter types, throwing a dynamic TypehintViolation-exception in case of a mismatch. In particular, the standard mocking framework used to expose an untyped API, using the catch-all mixed type. As a result, its use did not enjoy any static guarantees. Values flowing through mixed could encounter runtime enforcement boundaries, potentially leading to runtime crashes.
@@ 4.2. Problem Statement
Listing 6(a) presents an example diff, with unaltered lines for context and lines prefixed with + or - added or removed, respectively. In this diff, a developer improves the Application class (lines 16ff.) to be more object-oriented, introducing a new UserData class (lines 9–11), and altering existing code (lines 13 and 19) to make efficient use of it (line 14). However, the developer might not be aware of the already existing test (lines 1–7) for it, for example, because the test code might live in a completely different folder structure and is not easily discoverable.
In the case of the old, untyped mocking framework (Listing 6(b)), they would learn that they had broken a test only after CI ran, introducing significant delays: since the mocked method body now returns a dict value where UserData is expected, the Hack runtime throws a TypehintViolation. This leads to an additional cycle of checking out a diff, reproducing and debugging the unexpected failure, coming up with the right fix, and verifying it. The lack of static types for mocking has proven to be a hindrance to productivity, as indicated by the disproportionately high number of TypehintViolation-exceptions from CI logs.
@@
4.3. Methodology
Our experiments showed that we would be able to automatically migrate the majority (82%) of calls to the untyped mocking API onto the new typed variant (migratable call sites). The remainder of call sites (or unmigratable call sites) exercised the mocking API in unconventional ways and could not be auto-migrated.
We first migrated a random selection of half of the migratable call sites ($\sim$169,000 or 42% of all call sites in the codebase). During three weeks in 2024, we categorised $\sim$10,000 diffs:
  • Control Group (53.1% of diffs): Diffs that modified at least one unmigrated, but no partially or fully migrated files.
  • Test Group (24.5% of diffs): Diffs that modified at least one fully migrated, but no partially migrated or unmigrated files.
  • Mixed Group (22.4% of diffs): Diffs that modified at least one migrated and one unmigrated file, or at least one partially migrated file.
To estimate the effect of typed mocking, we employ a two-tailed Welch’s t-test, testing on the mean DAT difference between control and test groups. We discarded the mixed group as it contained diffs that touched both migrated and unmigrated files, making it ambiguous how to treat them. Because bigger diffs naturally require more coding time, we controlled for the diff size confounder using stratification.
@@ 4.4. Results
# of files changed % DAT saved $\frac{|Control|}{|Test|}$ pvaluep-value
1 $\sim 22\%$ $2.5$ <0.001<0.001
2 $\sim 22\%$ $2.3$ <0.001<0.001
@@
Variable $\sim 14\%$ $2.2$ <0.001<0.001
Table 1 summarizes the results of the stratified t-tests, revealing that use of the typed mocking API significantly reduced the time to write diffs, even after controlling for the number of files in each diff. The table shows the percentage of DAT saved, the relative control and test group sizes, and the p-value for each diff size category; diff size is defined as the number of files changed in a diff.
Diffs of all size categories showed statistically significant DAT reductions of approximately $14\%$ DAT ($p<0.001$, $n=\sim 10,000$). This was particularly prevalent for diff sizes up to four files, where diffs that used the typed mocking API showed an approximate 19–24% reduction in DAT over the control. We surmise that the inclusion of larger diffs (with more than four files changed), which might only tangentially touch mocking-related files, dilutes the larger savings seen in smaller diff size strata.
After seeing these positive results, Meta decided to fully migrate all 82% migratable call sites to typed mocking. While the results of this experiment have helped to quantify the direct effect of a Hack language feature and promoted a decision to roll it out across the code base, the wider decision-making processes in the organization also benefited. To the best of our knowledge, this is the first time a controlled, large-scale experiment demonstrated real-world productivity benefits of types in industry. Moreover, this result demonstrates that investing in low-latency static analysis tooling to shift correctness signals early in the development cycle can lead to substantial improvements in developer productivity, especially in large code bases.
@@ 5.1. Context
React is Meta’s JavaScript framework for developing dynamic user interfaces (Gackenheimer, 2015). Since its open-source release in 2013, it has become a standard for web and mobile apps (Inc., 2013). React components generate user interfaces based on the application’s current state. When the state changes, React updates the interface by re-evaluating components. This can be computationally intense. To optimize performance, developers often use memoization. In a sense, memoization is the front-end equivalent of caching. It prevents re-computing and re-drawing of unchanged components. This is especially important since components often form a deeply nested hierarchy and updating an unmemoized, but also unchanged parent component could cause a cascade of unnecessary re-renders.
@@ 5.2. Problem Statement
To avoid this, React developers can make their React components manually support memoization. This involves writing code to save component states manually; it has to be done for every single component. Manual memoization, similar to caching, not only causes a lot of non-trivial overhead unrelated to the actual feature engineers are implementing, but also introduces bugs and performance regressions, when done incorrectly (Savona and Zhang, 2023). For example, just to signal to the compiler that a certain component ExampleView supports memoization, a function’s signature had to look like this:
export default (React.memo<Props>(ExampleView): React.ComponentType<Props>);
instead of a simple
export default ExampleView;
The actual implementation logic inside the components was often dozens or hundreds of lines long.
For this reason, the React team set out to develop the next iteration of its bespoke React client-side compiler, React Forget. Its main feature is auto-memoization, which abstracts away manual memoization for the developer. With it, the compiler can determine the correct level and depth of memoization and export it automatically, eliminating the need for long export signatures and complex memoization logic.
@@ 5.3. Methodology and Results
Figure 7. Beanplots of DAT for diffs with manual (left) and automatic memoization (right) in a subset of React diffs.
While hard evidence regarding the usefulness of React Forget existed on operational metrics such as app performance and bug rates, it was initially impossible to evaluate its other purported aim, namely the impact on development productivity. With DAT, we could scientifically address it in a case study.
To obtain an initial, intuitive understanding of the difference the new React compiler might make to development velocity, we first compared diffs with auto-memoization to those without. Results ($n>>10,000$ diffs) in Figure 7 show that the DAT average, median, quartiles, and skew are significantly higher (i.e., worse) for diffs with manual memoization.
Based on these encouraging results, to obtain a precise difference between groups, and to take into account confounders, we built a mixed-effects regression model. This included lines of code, developer tenure, organization, time (modeled as month of the year), and others as its independent variables. It models $\Delta\mbox{DAT}$ as its output, to measure differences in mean DAT duration. As a post-hoc case study with non-randomized data, we use the Wasserstein distance to measure the difference between groups. It calculates the minimal transport distance between two probability distributions (Wikipedia, 2024) and is robust to noise and outliers, making it ideal for comparing distributions with different shapes or scales.
The final result of the mixed-effects model put $\Delta\mbox{DAT}$ between diffs with manual and auto-memoization at 33%. These DAT savings from the introduction of the React Forget compiler turned out greater than anticipated by domain experts.
@@ 6.1. Context
Meta’s development frameworks enable cross-platform code sharing, allowing code to be written once and deployed across multiple apps and platforms, reducing redundant efforts. One well-known example of such a framework is—React. Using code-sharing frameworks may save time compared to writing platform-specific code, because code only needs to be written once, but can be deployed to a number of different platforms. Developers report feeling more efficient with these frameworks, yet quantifying these gains was impossible before DAT.
@@ 6.2. Methodology
Counterfactual DAT aims to provide a realistic estimation of the development time that would have been required if code sharing had not been employed. While actual DAT for code-shared diffs is directly measurable, the counterfactual development time cannot be observed since these diffs were never created. Instead, we estimate it by identifying the platforms (such as Android) and apps (such as Instagram or Threads) targeted by each code-shared diff. For each platform, we compute a trimmed mean DAT from unshared diffs and normalize it by diff size, categorizing diffs into terciles based on lines of code modified.
The total counterfactual DAT is then the sum of the mean DAT for each targeted platform-app combination in the code-shared diff, controlled for diff size. Consequently, $DAT_{saved}$ can be negative if a code-shared diff incurs overhead, potentially due to the complexity of using or getting used to the shared libraries. Conversely, a higher count of targeted platform-app pairs increases the potential for actual DAT savings.
To illustrate this with a simple, hypothetical example, consider a code-shared diff developed using a framework that supports cross-platform deployment. This diff, with a total DAT of 1 hour, is deployed across four apps: Instagram Android, Instagram iOS, Facebook Android, and Facebook iOS. In contrast, implementing the same change independently for each app would require four separate diffs (this is the counterfactual to code sharing). Of course, we cannot simply multiply the actual DAT by this number to arrive at the counterfactual, but need to approximate it much closer, taking into account the idiosyncratic productivity in each different ecosystem. Querying the population diffs of comparable size might yield an average DAT of 0.9 hours for iOS and 0.6 hours for Android diffs (these are not the actual values). Therefore, the total counterfactual DAT is 3 hours (2 × 0.9 + 2 × 0.6), resulting in an estimated $DAT_{saved}$ of 2 hours.